# On the optimality of coin-betting for mean estimation

### v2

@@ Acknowledgements
Let $\mathcal{X}=\{0,1/2,1\}$, and $\mu=1/2$. Let $Q\in\mathcal{H}_{1/2}^{1}$ and let $q=Q(\{1/2\})$. The mean constraints reads $Q(\{1\})+q/2=1/2$, which implies that $Q(\{1\})=(1-q)/2$. The normalisation yields $Q(\{0\})=1-q-(1-q)/2=(1-q)/2$. This shows that there is a one-to-one correspondence between $[0,1]$ and $\mathcal{H}_{1/2}^{1}$. Now, since by definition there is a one-to-one correspondence between $\mathcal{H}_{1/2}^{1}$ and ^$\hat{\mathcal{H}}_{1/2}^{2}$, we obtain that we can parametrise ^$\hat{\mathcal{H}}_{1/2}^{2}$ by $[0,1]$. For $q\in[0,1]$, we hence let $P_{q}$ be the (unique) element of $\mathcal{H}_{1/2}^{2}$ that gives mass $q^{2}$ to $(1/2,1/2)$.

### v4

@@ A.2 Characterising the e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ with $\mathcal{X}=\{0,1/2,1\}$
Let $\mathcal{X}=\{0,1/2,1\}$, and $\mu=1/2$. Let $Q\in\mathcal{H}_{1/2}^{1}$ and let $q=Q(\{1/2\})$. The mean constraint reads $Q(\{1\})+q/2=1/2$, yielding $Q(\{1\})=(1-q)/2$. The normalisation yields $Q(\{0\})=1-q-(1-q)/2=(1-q)/2$. This shows that there is a one-to-one correspondence between [0,1][0,1][ 0 , 1 ] and $\mathcal{H}_{1/2}^{1}$. Now, since by definition there is a one-to-one correspondence between $\mathcal{H}_{1/2}^{1}$ and ^$\hat{\mathcal{H}}_{1/2}^{2}$, we obtain that we can parametrise ^$\hat{\mathcal{H}}_{1/2}^{2}$ by [0,1][0,1][ 0 , 1 ]. For $q\in[0,1]$, we hence let $P_{q}$ be the (unique) element of $\mathcal{H}_{1/2}^{2}$ that gives mass $q^{2}$ to (1/2,1/2)(1/2,1/2)( 1 / 2 , 1 / 2 ).